calculate the magnitude of the acceleration due to gravity on planet mercury mass = 2.99 x 10(power 23)kg and radius = 2.42 x 10(power 6) metres.
1
Expert's answer
2013-04-08T11:35:05-0400
Gravitational force is
&
F = G*m*M/r^2
&
where m is mass of body and M is mass of mercury, r is distance to the center (radius), G is gravitational constant.
Then the acceleration is
&
a=F/m =& G*M/r^2 = 6.67 x 10^(-11) x 2.99 x 10^(23) /( 2.42 x 10^6)^2 = 3.4 m/s^2

So, the magnitude of the acceleration due to gravity on planet mercury is 3.4 m/s^2